OpenStudy (anonymous):

The perimeter of the square equals the perimeter of the regular pentagon. Each side of the pentagon is 16 cm long. How long is each side of the square? I'll draw the thing

OpenStudy (anonymous):

 thats just one

OpenStudy (anonymous):

The pentagon has 5 sides. 5 * 16 = 80 The square must have the same perimeter of 80, but has only 4 sides 80 = 4 * x x = 20 The square has 20 cm sides

OpenStudy (anonymous):



OpenStudy (anonymous):

No pictures necessary, Bess :) The problem statement was very descriptive